Jan invested $1200 at a certain simple interest rate and $2200 at a rate of 3% higher. Her annual earnings were $253. Find the two interest rates if she earned $121 more on the larger investment than on the smaller.
what is the process and answer of this word problem? explain

Answer by lwsshak3(11628)

Jan invested $1200 at a certain simple interest rate and $2200 at a rate of 3% higher. Her annual earnings were $253. Find the two interest rates if she earned $121 more on the larger investment than on the smaller. What is the process and answer of this word problem? explain
**
let x=interest rate for $1200 investment
x+.03=interest rate for $2200 investment
..
2200(x+.03)+1200x=253
2200x+66+1200x=253
3400x=187
x=.055 or 5.5%
x+.03=.085 or 8.5%
Check:
Interest earned at $2200=.085*2200=187
Interest earned at $1200=.055*1200=66
difference=187-66=$121
..
ans:
Annual interest rate for larger $2200 investment = 8.5%
Annual interest rate for smaller $1200 investment = 5.5%